Consider a triatomic molecule, like CO2. The three atoms lie in a straight line, and we model the O atoms to be connected to the central C atom by a spring of constant k (same k for the two bonds). For this question, the atoms all oscillate in the plane of the molecule.

(a) What is the Lagrangian of the motion?

(b) What are the eigenfrequencies of the oscillations? (there are three, although one is zero)

(c) What are the ratios of the energies of the two non-zero frequencies?

(d) For the mode that oscillates with zero frequency, what is its motion? What about for the heighest frequency? It may help to find the eigenvectors.
Answer #1

Let a linear triatomic molecule Consists of tooo atoms 0, each of mass m,, placed at eithes side of thisd alom C The interactL= 출어 (역+)+노일을 *(912일을 . [T] K -K and [V - -K K 2k -K -k The Characteristicuati is k-t m k 2K-mk Km -k -K 0 k ave thyee ootexgies of the to0 fiequedcies is Ratios of mon- 20x0 K 2 R 1-9 K 2t2m, The motion fo oscillates the mode that ith Rexo fsegyDeterminatim eigenvectr: Let aigevecns case :1 W: 0, 0 be (,21M) Then, -K 0 -k 021 2k -k K -K katka 0 +Ka 1D =0 -ka t2ka21Ka312 K 2k-KW/ :0 22 -K -K 0 -Ka 22 3/M22-Ka20 a,2- M (Say) ano A2 2-2 yanvacta 2M1 2 Case 3: вз 1 eigen veca ) eiganvec given a- 2K к22 -0 - K KAA 013KA5 -ka23 -2K 0 V 12 Conditin TA[I] 0 2 O2M V Smpitinton On 7 M2 2.St Tim igenvectos he T 2 m, M2 + П2 / {0s} amd 2MThanks you.